complex analysis

2. Find the absolute values of - 2i(3 + 1)(2+41) (1 + 1) and and (3 + 4i)(-1 + 21) (-1 - 0) (3 - 1)
Answer #1

a liven -21(3 ti) (2+hi)liti) Now, -21 (3+1) (2 +41) (14i) = -2; (6+121+2i-4) Liti) --2162 +141)(2+i) =-2i(2+141+2i-14) =-28- -2713+i)(2+4) li+i)= 32+241 Absolute value of 2i3+ i)(2+41) (1Ai) = . Absolute value of 324241 = $132)²+(247² =√1024+526 =411-21 = (11-21) (4.2%) (4 +21)(4-21) 44-227-83-4 - 16+4 = 40-301 20 - ... (3+4i)(-+21) 4 3 Tunij (s-i) - Absolute value of (3